Need help on how to solve this...please include steps. Thank you :)
Assume that the real-world context can be accurately modeled with a linear function.
Between 1900 and 2003, the life expectancy for men and women in the United States increased dramatically. For men, life expectancy increased from 48 to 75 years. For women, life expectancy increased from 51 to 80 years.†
Model male life expectancy M as a function of female life expectancy f. (Round your coefficients to two decimal places.)

Answer by stanbon(75887)

Need help on how to solve this...please include steps. Thank you :)
Assume that the real-world context can be accurately modeled with a linear function.
Between 1900 and 2003, the life expectancy for men and women in the United States increased dramatically. For men, life expectancy increased from 48 to 75 years. For women, life expectancy increased from 51 to 80 years.†
Model male life expectancy M as a function of female life expectancy f. (Round your coefficients to two decimal places.)
------
In year 1800 you have ordered pair (51,45)
In year 2003 you have ordered pair (80,75)
---------
slope = (75-45)/(80-51) = 30/29
-----
Form:: y = mx + b
Solve for "b"::
75 = (30/29)*80 + b
b = -7.76
----
Equation: y = (30/29)x - 7.76
-----------
Cheers,
Stan H.
-----------